OpenStudy (anonymous):

A. Given the equation 5x − 4 = −2(3x + 2), solve for the variable. Explain each step and justify your process. B. Megan solved a similar equation below. Is Megan's solution correct? Explain why or why not. 3(2x − 4) = 5x − 1 6x − 12 = 5x − 1 11x − 12 = −1 11x = 1 x = 11

OpenStudy (anonymous):

@nincompoop

OpenStudy (hossamhoussien):

\[5x-4=-2(3x+2)\] \[5x-4=-6x-4\] \[5x=-6x-4+4\] \[5x=-6x\] \[5x+6x=0\] \[x=0\]

OpenStudy (hossamhoussien):

B) not true , you can solve it alone using the steps I did with part A if you still confused let me know

OpenStudy (anonymous):

given solution is not correct method to solve the equation 5x − 4 = −2(3x + 2) 5x-4=-6x-4 5x+6x-4=-4 11x=-4+4 11x=0 x=0 would be best explanation